OpenStudy (anonymous):

Lucas travels from home to work at a constant speed in 50 minutes. if he increases his speed by 25% over the course of the entire trip, how many minutes would it take for him to get to work

OpenStudy (anonymous):

Constant speed =x

OpenStudy (anonymous):

I am trying to figure out what's the initial speed

OpenStudy (anonymous):

distance=50 x (x+.25)=50x

OpenStudy (callisto):

\[s_{1} = d/t _{1} = d/50\]\[d = 50s _{1}\]\[s _{2} = 1.25 s _{1}\]\[s _{2} = d/t _{2}\]\[d = s _{2}t _{2}= 1.25 s _{1}t _{2}\]\[50s _{1}= s _{2}t _{2}=1.25 s _{1}t _{2}\]\[t _{2}=50s _{1}/ 1.25 s _{1} = 40 minutes\]

OpenStudy (callisto):

Hmm, it's a little bit messy..

OpenStudy (anonymous):

if goes 5/4 the speed, he gets there in 4/5 the time (40 mins).